# On Classes of Fredholm Type Operators

### 02 Jan 2024 v1

@@ 1 Introduction
In 1996, in [15, Definition 2.3], J.J. Koliha extended the notion of Drazin invertibility.
@@
Koliha [15] proved that $a\in A$ is generalized Drazin invertible if and only if there exists $\epsilon>0,$ such that for all $\lambda$ such that $0<\mid\lambda\mid<\epsilon,$ the element $a-\lambda e$ is invertible and he proved in [15, Theorem 4.2], that a generalized Drazin invertible element has a unique generalized Drazin inverse. He also proved that an element $a\in A$ is generalized Drazin invertible if and only if there exists an idempotent $p\in A$ commuting with $a,$ such that $a+p$ is invertible in $A$ and $ap$ is quasinilpotent.
@@ 3 On Pseudo B-Fredholm Operators
    It is proved in [5, Theorem 2.10] that an operator $T$ is a pseudo B-Fredholm operator if and only if $T$ is a compact perturbation of Riesz- Fredholm operator. In [9, Theorem 6.1], Boasso considered generalized Drazin invertible elements in the Calkin Algebra and he proved there the same result as [5, Theorem 2.10], though the proof of [5, Theorem 2.10] obtained later than [9, Theorem 6.1] was more direct. In fact the concept of p-invertibility enabled us to give left and right version of [9, Theorem 6.1].
  3. 3.
    The study of combination of operators of the type Riesz-Fredholm, Riesz-semi-Fredholm, Riesz-bounded below,Riesz-surjective,… appeared recently in several papers, see [12] and [18] for example and the references cited there.
@@
Consequently if $\lambda\in\mathbb{C}\setminus\{0\}$ and $\mid\lambda\mid$ is small enough, then $\Pi(T)-\lambda e$ is invertible. So if $0\in\sigma(\Pi(T)),$ then it is an isolated point. From [15, Theorem 3.1], it follows that $\Pi(T)$ is generalized Drazin invertible in $\mathcal{C}(X)$ and so $T$ is a pseudo B-Fredholm operator. ∎
@@
    In the recent works [2], [12] and [17], the authors studied operators similar to pseudo semi-B-Fredholm operators considered here. They defined them as the direct sum of a semi-Fredholm operator and a quasi-nilpotent one. However there exists operators which are pseudo semi-B-Fredholm operators in the sense of Definition 3.3, but do not have a decomposition as the direct sum of a semi-Fredholm operator and a quasi-nilpotent operator as seen by the following example.
  2. 2.
    The set of essentially left (right) generalized Drazin invertible operators was introduced and investigated in [13], An operator is essentially left (right)generalized Drazin invertible operator if and only if it is a direct sum of a quasinilpotent operator and a left (right) semi-Fredholm operator. Example 3.9 shows that the set of essentially left (right) generalized Drazin invertible operators which was introduced and investigated in [13], is contained strictly in the set of left (right) pseudo semi-B-Fredholm operators.
@@ 4 On Weak B-Fredholm Operators
1)2)1$\Rightarrow$ 2)1 ) ⇒ 2 ) Assume that $T$ is a weak B-Fredholm operator. So $\Pi(T)$ is Drazin invertible in $\mathcal{C}(X)$ and from [16, Proposition 4.9 ], there exists an idempotent $p\in\mathcal{C}(X)$ such that $\Pi(T)$ is $p$-invertible and $\Pi(T)p$ is nilpotent. So $\Pi(T)$ is left Drazin invertible and right Drazin invertible in $\mathcal{C}(X)$ and $T$ is a left and right weak semi-B-Fredholm operator.
2)1)2$\Rightarrow$ 1)2 ) ⇒ 1 ) Conversely assume now that $\Pi(T)$ is left and right Drazin invertible in $\mathcal{C}(X),$ so $\Pi(T)$ is left and right generalized Drazin invertible in $\mathcal{C}(X).$ From Theorem 3.7, it follows that $\Pi(T)$ is generalized Drazin invertible in $\mathcal{C}(X).$ Then from [15, Theorem 3.1] there exists an idempotent $p\in\mathcal{C}(X),$ commuting with $\Pi(T)$ such that $\Pi(T)p$ is quasi-nilpotent and $\Pi(T)+p$ is invertible. From [15, Theorem 3.1], we know that $p$ lies in the closed subalgebra of $\mathcal{C}(X)$ generated by $\Pi(T).$
@@
Remark 4.9.
The set of essentially left (right) Drazin invertible operators was introduced and investigated in [19], An operator is essentially left (right) Drazin invertible operator if and only if it is a direct sum of a nilpotent operator and a left (right) semi-Fredholm operator. Example 3.9 shows that the class of essentially left (right) Drazin invertible operators, is contained strictly in the class of left (right) weak semi-B-Fredholm operators.
Acknowledgment: The authors would like to thank Professor S. Č. Živković-Zlatanović for drawing their attention on the classes of essentially left (right) (generalized) Drazin invertible operators.

### 09 Jan 2024 v2

@@ 1 Introduction
In 1996, in [13, Definition 2.3], J.J. Koliha extended the notion of Drazin invertibility.
@@
Koliha [13] proved that $a\in A$ is generalized Drazin invertible if and only if there exists $\epsilon>0,$ such that for all $\lambda$ such that $0<\mid\lambda\mid<\epsilon,$ the element $a-\lambda e$ is invertible and he proved in [13, Theorem 4.2], that a generalized Drazin invertible element has a unique generalized Drazin inverse. He also proved that an element $a\in A$ is generalized Drazin invertible if and only if there exists an idempotent $p\in A$ commuting with $a,$ such that $a+p$ is invertible in $A$ and $ap$ is quasinilpotent.
@@ 3 On Pseudo B-Fredholm Operators
    In [9], in 2015, the author studied generalized Drazin invertible elements under Banach algebra homomorphisms. In the case of the Calkin Algebra, he proved in [9, Theorem 6.1] that an operator whose image is generalized Drazin invertible in the Calkin algebra is a Riesz-Fredholm operator. Later in 2019, in [5, Theorem 2.10], the authors proved the same result following using a shortened proof.
    Here in this paper, the concept of p-invertibility enable us to give left and right version of [9, Theorem 6.1].
@@
Consequently if $\lambda\in\mathbb{C}\setminus\{0\}$ and $\mid\lambda\mid$ is small enough, then $\Pi(T)-\lambda e$ is invertible. So if $0\in\sigma(\Pi(T)),$ then it is an isolated point. From [13, Theorem 3.1], it follows that $\Pi(T)$ is generalized Drazin invertible in $\mathcal{C}(X)$ and so $T$ is a pseudo B-Fredholm operator. ∎
@@
    In the recent works [2] and [15], the authors studied operators similar to pseudo semi-B-Fredholm operators considered here. They defined them as the direct sum of a semi-Fredholm operator and a quasi-nilpotent one. However there exists operators which are pseudo semi-B-Fredholm operators in the sense of Definition 3.3, but do not have a decomposition as the direct sum of a semi-Fredholm operator and a quasi-nilpotent operator as seen by the following example.
@@ 4 On Weak B-Fredholm Operators
1)2)1$\Rightarrow$ 2)1 ) ⇒ 2 ) Assume that $T$ is a weak B-Fredholm operator. So $\Pi(T)$ is Drazin invertible in $\mathcal{C}(X)$ and from [14, Proposition 4.9 ], there exists an idempotent $p\in\mathcal{C}(X)$ such that $\Pi(T)$ is $p$-invertible and $\Pi(T)p$ is nilpotent. So $\Pi(T)$ is left Drazin invertible and right Drazin invertible in $\mathcal{C}(X)$ and $T$ is a left and right weak semi-B-Fredholm operator.
2)1)2$\Rightarrow$ 1)2 ) ⇒ 1 ) Conversely assume now that $\Pi(T)$ is left and right Drazin invertible in $\mathcal{C}(X),$ so $\Pi(T)$ is left and right generalized Drazin invertible in $\mathcal{C}(X).$ From Theorem 3.7, it follows that $\Pi(T)$ is generalized Drazin invertible in $\mathcal{C}(X).$ Then from [13, Theorem 3.1] there exists an idempotent $p\in\mathcal{C}(X),$ commuting with $\Pi(T)$ such that $\Pi(T)p$ is quasi-nilpotent and $\Pi(T)+p$ is invertible. From [13, Theorem 3.1], we know that $p$ lies in the closed subalgebra of $\mathcal{C}(X)$ generated by $\Pi(T).$